OpenStudy (anonymous):

what is 16 divided by the square root of 6

OpenStudy (anonymous):

You might get better result if you post that in the math section ;)

OpenStudy (anonymous):



OpenStudy (kewlgeek555):

@magicstudy1903 WELCOME TO OPENSTUDY! :D Unfortunately, you asked you math question in the English section! Math section: http://openstudy.com/study#/groups/Mathematics Anyways...to answer your question. \[\frac{ 16 }{ \sqrt{6} }=\frac{ 8\sqrt{6} }{ 3 }\]